Circular No. 7801

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE B. Barris, University of Hawaii, on behalf of the High-Z Supernova

Search Team (cf. IAUC 6160, 6646, 7312, 7516, 7745, 7755, 7767, 7768), reports the discovery of more apparent supernovae on I-band images taken with the Subaru telescope (+ SUPRIME) by K. Chambers, M. Connelley, and D. Donovan; the new objects were identified by subtracting images taken on 2002 Jan. 9 from those taken on 2001 Oct. 22, Nov. 18, and 19. Spectral observations were made with the Keck I telescope (+ LRIS) by A. V. Filippenko and R. Chornock on Jan. 16 and 17 and with the Magellan telescope (+ LDSS2) by A. Clocchiatti on Jan. 17. The tabulated redshifts are of the host galaxy’s emission; magnitudes I_{1} and I_{2} are those on images from Jan. 9 and 14, respectively. Each apparent supernova is within 2′′ of its host galaxy’s center. None of the apparent supernovae below is present in images taken on 2001 Oct. 22 (limiting mag 26.0).

SN  2002 UT  α_{2000}  δ_{2000}  I_{1}  I_{2}  z
2002K Jan. 9 2^{h}26^{m}50.^{s}48 + 0^{o}43′30′′.1 24.0 24.2 0.644
2002L Jan. 9 2 26 55.69 + 0 21 37.8 23.8 24.6
2002M Jan. 9 2 27 13.14 + 0 44 41.2 23.4 23.6 0.379
2002N Jan. 9 2 27 51.91 + 0 28 53.6 23.2 23.7
2002O Jan. 9 2 28 22.17 + 0 49 46.5 23.7 24.0
2002P Jan. 9 2 29 05.71 + 0 47 20.1 23.8 23.2
2002Q Jan. 9 2 29 22.19 + 0 20 25.8 24.2 24.7
2002R Jan. 9 2 29 27.19 + 0 51 38.9 24.1 24.2
2002S Jan. 9 4 37 37.05 − 1 32 36.9 23.1 23.3
2002T Jan. 9 4 38 53.87 − 1 34 26.6 23.8 23.7
2002U Jan. 9 4 39 42.90 − 1 46 04.2 22.8 22.9
2002V Jan. 9 4 39 53.55 − 1 13 33.5 23.7 23.4
2002W Jan. 9 8 47 54.42 +44 13 42.9 24.0 23.7
2002X Jan. 9 8 48 30.54 +44 15 35.3 23.7 23.5 0.856

SUPERNOVA 2002F IN UGC 2885 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002F (cf. IAUC 7797), obtained by M. Calkins on Jan. 21.15 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova. There are incipient P-Cyg lines of hydrogen and helium on a blue continuum.

2002 January 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green